OpenStudy (anonymous):

How would I find the discriminant of a function using only a graph as my guide?

myininaya (myininaya):

like determining if is 0, negative, or positive or the actual number?

OpenStudy (anonymous):

You can tell whether the discriminant is positive, negative, or zero by how many times it crosses the x-axis. If it crosses twice, it is positive. If it touches the line once, it is zero. If it doesn't touch at all, there are no solutions and the discriminant will be negative.

myininaya (myininaya):

yes what jabber says

OpenStudy (anonymous):

Furthermore, and I'm just trying to figure this out here... if the function crosses twice, you should be able to determine the actual value of the descriminant based on the difference between the two 0s...

OpenStudy (anonymous):

Basically, the descriminant is b^2-4ac. Substituting D for discriminant, the difference between the two solutions of the polynomial is: (-b/2a + sqrt(D)/2a) - (-b/2a - sqrt(D)/2a) = sqrt(D)/2a +sqrt(D)/2a = 2sqrt(D)/2a = sqrt(D)/a To figure out what a is, check the y value when x is zero. So then, whatever the difference between the two roots of the function are, multiply that by a and then square it to get the descriminant. AM I RIGHT?

OpenStudy (anonymous):

Someone check my work please.